1. Express the sum in sigma notation: 12 + 22 + 32 +42 +52= 58-13
Answer #1

Givensumis B+2+3+4+5? we states the partial sums 2.S,=8 Sq=R+2= Sy=R+2+3? S, = 114 <+3 +4 Sg =R+2+3+4 +5? 班 个了 3 ーk K K-1 ~ K